OpenStudy (anonymous):

Solve sin2θ = 3cos2θ

OpenStudy (anonymous):

i believe the answer is pi/3+2pin,2pi/3+2pin,4pi/3+2pin,5pi/3+2pin

OpenStudy (anonymous):

Solve sin^2θ = 3cos^2θ

OpenStudy (freckles):

oh those are not double angles ?

OpenStudy (anonymous):

no squared, didnt realize the format change

OpenStudy (freckles):

\[\sin^2(\theta)=1-\cos^2(\theta) \] Therefore, \[1-\cos^2(\theta)=3\cos^2(\theta)\] get all the cos^2 's on one side and it will make more sense

OpenStudy (freckles):

\[1=4\cos^2(\theta) \\ \frac{1}{4}=\cos^2(\theta)\]

OpenStudy (freckles):

did you get to that equation I have there?

OpenStudy (anonymous):

which is 1.0471975511965977461542144610932, or pi/3. Because it is squared and negatives dont matter i should be able to add pi to that formula as many times as i want. I also should be able to subtract it from two pi. therefore the possible answers are pi/3+pin. and 5pi/3 + pin. However neither of those are options, only solutions involving two pi, so pi/3+2pin,2pi/3+2pin,4pi/3+2pin,5pi/3+2pin is the correct answer, right?

OpenStudy (freckles):

yeah \[\cos(\theta)=\frac{1}{2} \text{ gives us } \theta=\frac{\pi}{3} +2npi , \theta=\frac{5\pi}{3}+2npi \\ \cos(\theta)=\frac{-1}{2} \text{ gives us } \theta=\frac{2\pi}{3}+2n \pi, \theta=\frac{4\pi}{3}+2npi\] this can be found on the unit circle though

OpenStudy (anonymous):

Ok thanks:) yeah the subtract from two pi thing is what i meant for "found on the unit circle"

OpenStudy (freckles):

like we just look for when the x coordinate is 1/2 for the first one and for the second one just look for when the x coordinate is -1/2